OpenStudy (anonymous):

What are the vertices of the hyperbola given by the equation 16x2 - y2 + 96x + 10y + 103 = 0?

OpenStudy (anonymous):

16x² – y² + 96x + 10y + 103 = 0 You need to put the equation into standard hyperbola form. regroup terms 16x² + 96x – y² + 10y = -103 factor out leading coefficients 16(x² + 6x) – (y² - 10y) = -103 complete the squares 16(x² + 6x + 3²) – (y² - 10y + (-5)²) = -103 +16(3²) - (-5)² 16(x + 3)² – (y - 5)² = 16 divide by 16 to set righthand-side to 1 16(x + 3)²/16 – (y - 5)²/16 = 16/16 (x + 3)² – (y - 5)²/16 = 1 express denominators as squares (x + 3)²/1² – (y - 5)²/4² = 1 center at (-3, 5) The x term is positive, so: the hyperbola is horizontal. vertices (-3±1, 5) co-vertices (-3, 5±4)